You and two friends are going to rent a
condo downtown
The condo has 3 bedrooms
Bedroom #1 11 x 11
Bedroom #2 12 X 14
Bedroom #3 15 X 15
The monthly rental for the condo is $4,500
how much each person should pay
Show your work and calculate your answer.
Answer #1

Answer:

a) No. of persons = 3

b) No. of bedrooms of the condo = 3

c) Size of bedroom #1 = 11' X 11' = 121 square feet.

d) Size of bedroom #2 = 12' X 14' = 168 square feet.

e) Size of bedroom #3 = 15' X 15' = 225 square feet.

f) Hence, total size of 3 bedroom = 121 square feet + 168 square feet + 225 square feet = 514 square feet.

g) The monthly rental for the condo = $4,500

h) Hence, for 514 square feet monthly rental = $4,500

i) Hence, the monthly rental for bedroom #1 = $4,500/514 X 121 = $1,059 (Rounded)

j) Hence, the monthly rental for bedroom #2 = $4,500/514 X 168 = $1,471 (Rounded)

k) Hence, the monthly rental for bedroom #3 = $4,500/514 X 225 = $1,970 (Rounded)

So, from the above it can concluded that as per occupancy of the room each should pay $1,059, $1,471 & $1,970 for bedroom #1,2 & 3 respectively.